# MyShadeQuoteStyleleftmargin=0pt, rightmargin=0pt, backgroundcolor=gray!25, linewidth=0pt, skipbelow=skipabove=

### v2

@@ 6.2. Threats to Validity
Internal Validity A common threat to the validity of our study is our hyper-parameter settings for DLVD models and LLMs. For DLVD, hyper-parameter tuning is extremely expensive given the scale of our study, hence we followed the setting of previous studies (Nong et al., 2024, 2023). In addition, when comparing different data augmentation approaches we used the same settings for DLVD models, which ensures our comparisons are fair. For LLMs, we mostly used the defaults, and only set the number of new tokens to ChatGPT’s maximum, and lowered ChatGPT’s temperature to 0.5. Better results might be achievable by fine-tuning these hyper-parameters. Another threat is that LLMs might not always follow the prompts as instructed and could produce hallucinations. To alleviate this threat, we add a verifier component to filter out low-quality code, and we use a retry mechanism that prompts the LLM up to three times in case of errors (including an empty code response). More importantly, our goal in this study is to augment existing vulnerable datasets to help models capture the vulnerability patterns and generalize better, rather than generating high-quality vulnerable samples. Completely correct code is not necessary for training models (Wang et al., 2024) and data with subtle noise helps with the generalization of models (Bishop, 1995; Frenay and Verleysen, 2014; Hickey, 1996). We acknowledge that more sophisticated methods can be used for the verification phase, but starting from a simple module (i.e. a parser) was necessary for evaluating whether using LLMs for vulnerability detection is even feasible.

### v3

@@ 6.3. Cost Analysis
6.4. Quality Analysis of Augmented Vulnerable Samples
We focus on data augmentation and not generation in this study, and introducing noise is inevitable. This is common in the application of other data augmentation techniques in the fields of natural language processing and computer vision (Zhang et al., 2018; Guo et al., 2019; DeVries and Taylor, 2017; Olsson et al., 2021; Yun et al., 2019). It is crucial to realize that data augmentation is used as a means to improve the generalizability of the models. A reasonable amount of noise can help with the regularization of the model. Otherwise, performance degradation will be seen. In addition, based on the results of recent studies (Li et al., 2022; Dong et al., 2025) show that code data augmentation techniques that slightly break the syntax can help the training of the models. However, we recognize the importance of evaluating the quality of augmented vulnerable samples to provide practical insights for practitioners when using our approach.
To this end, we conducted a manual assessment of the generated code. Specifically, we randomly selected 50 samples per augmentation strategy (i.e., Mutation, Injection, and Extension), produced using ChatGPT-3.5 in RQ1 (150 samples in total). We manually inspected each sample to determine whether the LLM followed our instructions and successfully generated a vulnerable example that preserved the intended vulnerability pattern. Our analysis revealed that 72%, 82%, and 90% of the samples generated via Mutation, Injection, and Extension, respectively, were correctly implemented and truly vulnerable. These results indicate an acceptable level of quality. The higher success rates for Injection and Extension may help explain their superior performance compared to Mutation. This also suggests that LLMs may find it easier to perform injection and extension tasks than mutation. We also study the failure cases, and we observe certain patterns. For instance, the common case for the Mutation was changing the code in a way different from the methods mentioned, such that the output is significantly different from the instruction. For Injection, wrong injections or half-complete injections were more common. For Extension, this had to do with the removal of the input code instead of extending it.
@@ 6.5. Threats to Validity
Internal Validity A common threat to the validity of our study is our hyperparameter settings for DLVD models and LLMs. For DLVD, hyperparameter tuning is extremely expensive given the scale of our study; as such, we followed the setting of previous studies (Nong et al., 2024, 2023). In addition, when comparing different data augmentation approaches, we used the same settings for DLVD models, which ensures our comparisons are fair. For LLMs, we mostly used the defaults, and only set the number of new tokens to ChatGPT’s maximum, and lowered ChatGPT’s temperature to 0.5. Better results might be achievable by fine-tuning these hyperparameters.
Label noise in the datasets can pose another threat to the validity of our experiments. Particularly, Devign and BigVul, even though they are commonly used in previous studies, are known to be noisy datasets (Croft et al., 2023; Risse and Böhme, 2024; Ding et al., 2024), meaning that a considerable number of vulnerable samples are incorrectly labeled, which can affect the results of our experiments. To ensure these results can be trusted, we expanded our study by dedicating an additional RQ to using PrimeVul, which is a more recent dataset with considerably less label noise, and showed that our results still hold.
Another threat is that LLMs might not always follow the prompts as instructed and could produce hallucinations. To alleviate this threat, we add a verifier component to filter out low-quality code, and we use a retry mechanism that prompts the LLM up to three times in case of errors (including an empty code response). In addition, we study the quality of augmented vulnerable samples, and found that the quality is at an acceptable level. We encourage future research to integrate more advanced approaches to reduce hallucinations. More importantly, our goal in this study is to augment existing vulnerable datasets to help models capture the vulnerability patterns and generalize better, rather than generating high-quality vulnerable samples. Slightly breaking the syntax and not having completely correct code remains beneficial for training of the models (Wang et al., 2024; Li et al., 2022; Dong et al., 2025), and data with subtle noise helps with the generalization of models (Bishop, 1995; Frenay and Verleysen, 2014; Hickey, 1996). We acknowledge that more sophisticated methods can be used for the verification phase, but starting from a simple module (i.e., a parser) was necessary for evaluating whether using LLMs for vulnerability detection is even feasible.